OpenStudy (anonymous):

log(x) + log(x - 1) = log(4x)

OpenStudy (anonymous):

start with \[\log(x(x-1))=\log(4x)\] then solve \[x(x-1)=4x\] for \(x\) don't forget to check your answers as one might make the input negative (forbidden with logs)